The accounting record for Katzen Company reported the following selected information:

Operating Expenses

$270,000

Sales Returns and Allowances

78,000

Sales Discounts

36,000

Sales Revenue

1,050,000

Cost of Goods Sold

402,000


Determine Katzen Company's gross profit.

Select one:

A. $462,000

B. $498,000

C. $534,000

D. $420,000
Answer #1

Net sales=Sales Revenue-Sales Returns and Allowances-Sales Discounts

=1,050,000-78,000-36,000=$936000

Gross profit=Net sales-Cost of Goods Sold

=936000-402,000

=$534,000